# Asymptotic error distribution for stochastic Runge–Kutta methods of strong order one

### v1

@@ 3.2. An appurtenant numerical method
For $a=(a_{1},\ldots,a_{d})^{\top},~{}b=(b_{1},\ldots,b_{d})^{\top}\in\mathbb{R}^{d}$, denote $a*b\in\mathbb{R}^{d}$ with $(a*b)_{i}=a_{i}b_{i}$, $i=1,\ldots,d$ and $a^{m}\in\mathbb{R}^{d}$ with $(a^{m})_{i}=a_{i}^{m}$. The following lemma gives a sufficient expansion of $Y_{t,y}^{SRK}(t+h)$, which helps us to find the appurtenant numerical method as in Framework 1.
@@
and for any $p\geq 1$, $\|R^{2}_{Y^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\alpha^{\prime}})h^{5/2}$ for some $\alpha^{\prime}\geq 1$.
Proof.
By (3.20)-(3.21), we can write
$\;+\frac{1}{2}\Delta\widehat{W}^{2}(\sum_{j=1}^{s_{0}}b_{ij})^{2}\mathcal{D}^{2}g(y)(g(y),g(y))+\bar{R}^{3}_{g,i},,$ (3.26)
with $\|\bar{R}^{3}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{3/2}$. Plugging (3.18) and (3.26) into (3.14) produces
$\;+\Delta\widehat{W}^{3}\sum_{j,k,l=1}^{s_{0}}b_{ij}b_{jk}b_{kl}(\nabla g(y))^{2}g(y)+\frac{1}{2}\Delta\widehat{W}^{3}\sum_{j=1}^{s_{0}}b_{ij}\big{(}\sum_{k=1}^{s_{0}}b_{jk}\big{)}^{2}\mathcal{D}^{2}g(y)(g(y),g(y))+R^{3}_{Z_{i}},$ (3.27)
The Taylor formula leads to
$\|\tilde{R}^{3}_{f,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{3/2}.$ (3.28)
Plugging (3.19) into $\nabla f(y)(Z_{i}-y)$ and (3.15) into $\frac{1}{2}\mathcal{D}^{2}f(y)(Z_{i}-y,Z_{i}-y)$ yields
$\;+\Delta\widehat{W}^{2}\sum_{j=1}^{s_{0}}b_{ij}\sum_{k=1}^{s_{0}}b_{jk}\nabla f(y)\nabla g(y)g(y)+\frac{1}{2}\Delta\widehat{W}^{2}(\sum_{j=1}^{s_{0}}b_{ij})^{2}\mathcal{D}^{2}f(y)(g(y),g(y))+R^{3}_{f,i},$ (3.29)
where $R^{3}_{f,i}:=\nabla f(y)R^{2}_{Z_{i}}+\Delta\widehat{W}\sum_{j=1}^{s_{0}}b_{ij}\mathcal{D}^{2}f(y)(g(y),R^{1}_{Z_{i}})+\frac{1}{2}\mathcal{D}^{2}f(y)(R^{1}_{Z_{i}},R^{1}_{Z_{i}})+\tilde{R}^{3}_{f,i}$. Combining the estimates for $R^{1}_{Z_{i}}$ and $R^{2}_{Z_{i}}$ from the proof of Theorem 3.3 and using (3.28), one has $\|R^{3}_{f,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{3/2}.$
Applying the Taylor formula to $g(Z_{i})$, one has
where $\tilde{R}^{4}_{g,i}:=\frac{1}{6}\int_{0}^{1}(1-\lambda)^{3}\mathcal{D}^{4}g(y+\lambda(Z_{i}-y))(Z_{i}-y,Z_{i}-y,Z_{i}-y,Z_{i}-y)\mathrm{d}\lambda$. By $\mathcal{D}^{4}g\in\mathbf{F}$ and Lemma 3.1, $\|\tilde{R}^{4}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{2}$. Plugging (3.27), (3.19) and (3.15) into $\nabla g(y)(Z_{i}-y)$, $\mathcal{D}^{2}g(y)(Z_{i}-y,Z_{i}-y)$ and $\mathcal{D}^{3}g(y)(Z_{i}-y,Z_{i}-y,Z_{i}-y)$, respectively, and keeping those terms with strong order no larger than $1.5$, we obtain
$\;+\sum_{j=1}^{s_{0}}b_{ij}\sum_{k,l=1}^{s_{0}}b_{ik}b_{kl}\mathcal{D}^{2}g(y)(g(y),\nabla g(y)g(y))+\frac{1}{6}(\sum_{j=1}^{s_{0}}b_{ij})^{3}\mathcal{D}^{3}g(y)(g(y),g(y),g(y))\Big{]}+R^{4}_{g,i}$ (3.30)
with $\|R^{4}_{g,i}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\iota})h^{2}$.
Plugging (3.2)-(3.30) into (3.10) and using $\alpha^{\top}e=\beta^{\top}e=1$ and $\beta^{\top}(Be)=\frac{1}{2}$, we obtain (3.25) with $R^{2}_{Y^{SRK}}=h\sum_{i=1}^{s_{0}}\alpha_{i}R^{3}_{f,i}+\Delta\widehat{W}\sum_{i=1}^{s_{0}}\beta_{i}R^{4}_{g,i}$. Finally, $\|R^{2}_{Y^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\alpha^{\prime}})h^{5/2}$ is obtained due to the previous estimates for $R^{3}_{f,i}$ and $R^{4}_{g,i}$. ∎
@@
Let $\{\bar{Y}_{n}\}_{n=0}^{N}$ be the numerical solution generated by $\bar{Y}_{t,h}(t+h)$, i.e., $\bar{Y}_{n+1}=\bar{Y}_{t_{n},\bar{Y}_{n}}(t_{n+1})$, $n=0,\ldots,N-1$ and $\bar{Y}_{0}=Y_{0}$. Based on Lemma 3.5 and Theorem 2.1, it can be shown that the mean-square error order between $\bar{Y}_{n}-Y^{SRK}_{n}$ is $1.5$, and $\{\bar{Y}_{n}\}_{n=0}^{N}$ has mean-square order $1$ when approximating (3.1). This is to say, $\{\bar{Y}_{n}\}_{n=0}^{N}$ is a candidate of the appurtenant method of the SRK method (3.9). However, the terms $\Delta W^{i}$, $i=2,3,4$ in (3.2) is very inconvenient for us to give a proper continuous version of $\{\bar{Y}_{n}\}_{n=0}^{N}$ for studying its asymptotic error distribution. For this end, we use the multiple Itô integral to represent $\Delta W^{i}$, $i=2,3,4$, and give the appurtenant method of (3.9).
Noting that $(W(s)-W(t))^{2}=2\int_{t}^{s}(W(r)-W(t))\mathrm{d}W(r)+(s-t)$, we have
@@
due to [17, Page 102, Lemma 2.2]. This combined with Lemma 3.2 verifies the condition (A2) of Theorem 2.1.
@@
By Theorem 3.3 and Lemma 3.6, we have that $\widetilde{Y}_{n}$ converges to $Y(t_{n})$ of mean-square order $1$. Thus, $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ is an appurtenant method of (3.9). In fact, we can prove that $\sup_{n=0,1,\ldots,N}\|\widetilde{Y}_{n}-Y(t_{n})\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)h$ for any $p\geq 2$.
@@
It follows from (3.24) and (3.32) that
@@
Comparing $Y_{t,y}(t+h)$ and $\widetilde{Y}_{t,y}(t+h)$, one can prove $Y_{t,y}(t+h)-\widetilde{Y}_{t,y}(t+h)$ satisfies the condition (A1) of Theorem 2.1 for any $p\geq 1$ with $p_{1}=2$ and $p_{2}=\frac{3}{2}$. In addition, the condition (A2) of Theorem 2.1 holds due to (3.3) and (3.38). Also, $Y_{t,y}(t+h)$ fulfills the condition (A3) of Theorem 2.1; see [22, Lemma 2.2]. Thus, the proof is complete based on Theorem 2.1. ∎
@@
Let Assumption 2 hold. Then $N(Y^{SRK}_{N}-Y(T))$ and $N(\widetilde{Y}_{N}-Y(T))$ has the same limit distribution as $N\to\infty$, if one of them converges in distribution.
@@
Note that $N(Y^{SRK}_{N}-Y(T))=N(\widetilde{Y}_{N}-Y(T))+Q_{N}$ with $Q_{N}=N(Y^{SRK}_{N}-\widetilde{Y}_{N})$. By Lemma 3.6, $Q_{N}$ converges to 00 in probability. Thus, the proof is finished as a result of Slutzky’s theorem. ∎
@@ 4.1. Limit distribution of normalized error process
where $\kappa_{N}(s)=\lfloor\frac{s}{h}\rfloor h=\lfloor\frac{Ns}{T}\rfloor\frac{T}{N}$. Then one has $\widetilde{Y}(t_{n})=\widetilde{Y}_{n}$, $n=0,\ldots,N$ from (3.37) and (4.1). It is not difficult to show that $\widetilde{Y}$ converges to $Y$ with strong order $1$.
@@
We are in the position to give a expansion for the normalized error $N(\widetilde{Y}(t)-Y(t))$.
@@
$\;-\frac{N}{2}\int_{0}^{t}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))(s-\kappa_{N}(s))\mathrm{d}s+\tilde{R}_{3}^{N}(t).$ (4.8)
@@
Let Assumption 2 hold. Then for any $t\in[0,T]$, $\mathbf{E}|U^{N}(t)-\widetilde{U}^{N}(t)|^{2}=0$. Here, $\widetilde{U}^{N}$ is the strong solution of the following equation
@@
$\lim_{N\to\infty}\mathbf{E}\Big{|}\int_{0}^{t}\big{(}H(\widetilde{Y}(\kappa_{N}(s)))-H(Y(s))\Big{)}\mathrm{d}s\Big{|}^{2}$ $=0,$ (4.14)
$\lim_{N\to\infty}\mathbf{E}\Big{|}\int_{0}^{t}\big{(}H(\widetilde{Y}(\kappa_{N}(s)))-H(Y(s))\Big{)}\mathrm{d}W(s)\Big{|}^{2}$ $=0.$
Thus, it holds that $\lim\limits_{N\to\infty}\mathbf{E}|I^{N}_{i}(t)-I_{i}(t)|^{2}=0$ for $i=1,2$ and $t\in[0,T]$. Moreover, applying (4.14) and [5, Proposition 4.2], we obtain $\lim\limits_{N\to\infty}\mathbf{E}|I^{N}_{3}(t)-I_{3}(t)|^{2}=0$ for any $t\in[0,T]$.
@@
[…]
@@
The proof is mainly based on [10, Theorem 4-1]. Hereafter, denote by $\langle X,Y\rangle_{t}$, $t\in[0,T]$ the cross variation process between the real-valued semi-martingales $\{X(t)\}_{t\in[0,T]}$ and $\{Y(t)\}_{t\in[0,T]}$. Let $I_{4}^{N,i}$ denote the $i$th component of $I_{4}^{N}$ for $i=1,\ldots,d$. Recall that
@@
Applying (4.14) and [5, Proposition 4.2], we arrive at $\lim\limits_{N\to\infty}\mathbf{E}|B_{1}^{N}(t)-\frac{T}{2}\int_{0}^{t}G_{1}^{i}(Y(s))\mathrm{d}s|^{2}=0$. Further, similar to the proof of (4.9), one has $\lim\limits_{N\to\infty}\mathbf{E}|B_{2}^{N}(t)|^{2}=0$. Thus, for any $t\in[0,T]$ and $i=1,\ldots,d$,
@@
Similar to the proof of (4.9), one can show that $\mathbf{E}|B_{3}^{N}(t)|^{2}+\mathbf{E}|B_{4}^{N}(t)|^{2}\leq Kh\to 0$, as $N\to\infty$. Further, using [5, Proposition 4.2] yields $\int_{0}^{t}G_{l}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{l}^{j}(\widetilde{Y}(\kappa_{N}(s)))\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}^{2}\mathrm{d}s$ converges to $\frac{1}{3}\int_{0}^{t}G_{l}^{i}(Y(s))G_{l}^{j}(Y(s))\mathrm{d}s$ in $\mathbf{L}^{2}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$ for $l=1,2$. In this way, we obtain
@@
[…]
@@
Further, it follows from [10, Proposition 1-4] that $\Psi^{i}$, $i=1,\ldots,d$ can be represented as
@@
$\Phi(t)=\int_{0}^{t}\nabla\bar{f}(Y(s))\Phi(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))\Phi(s)\mathrm{d}W(s)+\sum_{i=1}^{4}I_{i}(t).$

### v2

@@ 3.2. An appurtenant numerical method
For $a=(a_{1},\ldots,a_{d})^{\top},~b=(b_{1},\ldots,b_{d})^{\top}\in\mathbb{R}^{d}$, denote $a*b\in\mathbb{R}^{d}$ with $(a*b)_{i}=a_{i}b_{i}$, $i=1,\ldots,d$ and $a^{m}\in\mathbb{R}^{d}$ with $(a^{m})_{i}=a_{i}^{m}$. The following lemma gives a sufficient expansion of $Y_{t,y}^{SRK}(t+h)$, helping us to find the appurtenant numerical method as in Framework 1, whose proof is deferred to the appendix.
@@
and there is $\alpha^{\prime}\geq 1$ such that for any $p\geq 1$, $\|R^{2}_{Y^{SRK}}\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)(1+|y|^{\alpha^{\prime}})h^{5/2}$.
@@
Let $\{\bar{Y}_{n}\}_{n=0}^{N}$ be the numerical solution generated by $\bar{Y}_{t,h}(t+h)$, i.e., $\bar{Y}_{n+1}=\bar{Y}_{t_{n},\bar{Y}_{n}}(t_{n+1})$, $n=0,\ldots,N-1$ and $\bar{Y}_{0}=Y_{0}$. Based on Lemma 3.5 and Theorem 2.1, it can be shown that the mean-square error order between $\bar{Y}_{n}-Y^{SRK}_{n}$ is 1.51.51.5, and $\{\bar{Y}_{n}\}_{n=0}^{N}$ has mean-square order 111 when approximating (3.1). This is to say, $\{\bar{Y}_{n}\}_{n=0}^{N}$ is a candidate of the appurtenant method of the SRK method (3.9). However, the terms $\Delta W^{i}$, $i=2,3,4$ in (3.2) are very inconvenient for us to give a proper continuous version of $\{\bar{Y}_{n}\}_{n=0}^{N}$ for studying its asymptotic error distribution. For this end, we use the multiple Itô integral to represent $\Delta W^{i}$, $i=2,3,4$, and give the appurtenant method of (3.9).
First, it holds that
@@
due to [16, Page 102, Lemma 2.2]. This combined with Lemma 3.2 verifies the condition (A2) of Theorem 2.1.
@@
By Theorem 3.3 and Lemma 3.6, we have that $\widetilde{Y}_{n}$ converges to $Y(t_{n})$ of mean-square order 111. Thus, $\{\widetilde{Y}_{n}\}_{n=0}^{N}$ is an appurtenant method of (3.9). In fact, we can prove that $\sup_{n=0,1,\ldots,N}\|\widetilde{Y}_{n}-Y(t_{n})\|_{\mathbf{L}^{p}(\Omega)}\leq K(p)h$ for any $p\geq 2$.
@@
It follows from (A.13) and (3.14) that
@@
Comparing $Y_{t,y}(t+h)$ and $\widetilde{Y}_{t,y}(t+h)$, one can prove $Y_{t,y}(t+h)-\widetilde{Y}_{t,y}(t+h)$ satisfies the condition (A1) of Theorem 2.1 for any $p\geq 1$ with $p_{1}=2$ and $p_{2}=\frac{3}{2}$. In addition, the condition (A2) of Theorem 2.1 holds due to (3.3) and (3.20). Also, $Y_{t,y}(t+h)$ fulfills the condition (A3) of Theorem 2.1; see [21, Lemma 2.2]. Thus, the proof is complete based on Theorem 2.1. ∎
@@
Let Assumption 2 hold. Then $N(Y^{SRK}_{N}-Y(T))$ and $N(\widetilde{Y}_{N}-Y(T))$ have the same limit distribution as $N\to\infty$, if one of them converges in distribution.
@@
Note that $N(Y^{SRK}_{N}-Y(T))=N(\widetilde{Y}_{N}-Y(T))+Q_{N}$ with $Q_{N}=N(Y^{SRK}_{N}-\widetilde{Y}_{N})$. By Lemma 3.6, $Q_{N}$ converges to 0 in probability. Thus, the proof is finished as a result of Slutzky’s theorem. ∎
@@ 4.1. Limit distribution of normalized error process
where $\kappa_{N}(s)=\lfloor\frac{s}{h}\rfloor h=\lfloor\frac{Ns}{T}\rfloor\frac{T}{N}$. Then one has $\widetilde{Y}(t_{n})=\widetilde{Y}_{n}$, $n=0,\ldots,N$ from (3.19) and (4.1). It is not difficult to show that $\widetilde{Y}$ converges to $Y$ with strong order 111.
@@
We are in the position to give an expansion for the normalized error $N(\widetilde{Y}(t)-Y(t))$.
@@
$\;-\frac{N}{2}\int_{0}^{t}\mathcal{D}^{2}\bar{f}(\widetilde{Y}(\kappa_{N}(s)))(g(\widetilde{Y}(\kappa_{N}(s))),g(\widetilde{Y}(\kappa_{N}(s))))(s-\kappa_{N}(s))\mathrm{d}s+\tilde{R}_{3}^{N}(t),$ (4.8)
@@
Let Assumption 2 hold. Then for any $t\in[0,T]$, $\lim\limits_{N\to\infty}\mathbf{E}|U^{N}(t)-\widetilde{U}^{N}(t)|^{2}=0$. Here, $\widetilde{U}^{N}$ is the strong solution of the following equation
@@
$\lim_{N\to\infty}\mathbf{E}\Big{|}\int_{0}^{t}\big{(}H(\widetilde{Y}(\kappa_{N}(s)))-H(Y(s))\Big{)}\mathrm{d}s\Big{|}^{2}$ =0=0,= 0 , (4.14)
$\lim_{N\to\infty}\mathbf{E}\Big{|}\int_{0}^{t}\big{(}H(\widetilde{Y}(\kappa_{N}(s)))-H(Y(s))\Big{)}\mathrm{d}W(s)\Big{|}^{2}$ =0=0.= 0 .
Thus, it holds that $\lim\limits_{N\to\infty}\mathbf{E}|I^{N}_{i}(t)-I_{i}(t)|^{2}=0$ for $i=1,2$ and $t\in[0,T]$. Moreover, applying (4.14) and [6, Proposition 4.2], we obtain $\lim\limits_{N\to\infty}\mathbf{E}|I^{N}_{3}(t)-I_{3}(t)|^{2}=0$ for any $t\in[0,T]$.
@@
Let Assumption 2 hold. Then $I_{4}^{N}\overset{satbly}{\Longrightarrow}I_{4}$ as $\mathbf{C}([0,T];\mathbb{R}^{d})$-valued random variables as $N\to\infty$. Here, $I_{4}$ is the stochastic process defined by
@@
The proof is mainly based on [9, Theorem 4-1]. Hereafter, denote by $\langle X,Y\rangle_{t}$, $t\in[0,T]$ the cross variation process between the real-valued semi-martingales $\{X(t)\}_{t\in[0,T]}$ and $\{Y(t)\}_{t\in[0,T]}$. Let $I_{4}^{N,i}$ denote the $i$th component of $I_{4}^{N}$ for $i=1,\ldots,d$. Recall that
@@
Applying (4.14) and [6, Proposition 4.2], we arrive at $\lim\limits_{N\to\infty}\mathbf{E}|B_{1}^{N}(t)-\frac{T}{2}\int_{0}^{t}G_{1}^{i}(Y(s))\mathrm{d}s|^{2}=0$. Further, similar to the proof of (4.9), one has $\lim\limits_{N\to\infty}\mathbf{E}|B_{2}^{N}(t)|^{2}=0$. Thus, for any $t\in[0,T]$ and $i=1,\ldots,d$,
@@
Similar to the proof of (4.9), one can show that $\mathbf{E}|B_{3}^{N}(t)|^{2}+\mathbf{E}|B_{4}^{N}(t)|^{2}\leq Kh\to 0$, as $N\to\infty$. Further, using [6, Proposition 4.2] yields $\int_{0}^{t}G_{l}^{i}(\widetilde{Y}(\kappa_{N}(s)))G_{l}^{j}(\widetilde{Y}(\kappa_{N}(s)))\big{(}\frac{Ns}{T}-\lfloor\frac{Ns}{T}\rfloor\big{)}^{2}\mathrm{d}s$ converges to $\frac{1}{3}\int_{0}^{t}G_{l}^{i}(Y(s))G_{l}^{j}(Y(s))\mathrm{d}s$ in $\mathbf{L}^{2}(\Omega,\mathcal{F},\mathbf{P};\mathbb{R}^{d})$ for $l=1,2$. In this way, we obtain
@@
Combining (4.15)-(4.16) and using [9, Theorem 4-1], we obtain $I_{4}^{N}\overset{stably}{\Longrightarrow}\Psi$ as $\mathbf{C}([0,T];\mathbb{R}^{d})$-valued random variables as $N\to\infty$, and
@@
Further, it follows from [9, Proposition 1-4] that $\Psi^{i}$, $i=1,\ldots,d$ can be represented as
@@
By Proposition 2.2 and Lemma 4.4, $(I_{1},I_{2},I_{3},I_{4}^{N})\overset{stably}{\Longrightarrow}(I_{1},I_{2},I_{3},I_{4})$ in $\mathbf{C}([0,T];\mathbb{R}^{d})^{\otimes 4}$ as $N\to\infty$. Note that a continuous mapping maps a family of random variables converging stably in law into another family of random variables converging stably in law, which can be directly shown according to the definition of stable convergence. Thus, we obtain $\Phi^{N}:=I_{1}+I_{2}+I_{3}+I_{4}^{N}\overset{stably}{\Longrightarrow}\Phi:=I_{1}+I_{2}+I_{3}+I_{4}$ in $\mathbf{C}([0,T];\mathbb{R}^{d})$.
Note that the coefficients $\nabla\bar{f}$ and $\nabla g$ of (4.12) are bounded and $\mathcal{D}^{2}\bar{f},\mathcal{D}^{2}g\in\mathbf{F}$. In addition, a direct computation leads to
$\sup_{N\geq 1}\sup_{t\in[0,T]}\mathbf{E}|\Phi^{N}(t)|^{4}+\sup_{t\in[0,T]}\mathbf{E}|\Phi(t)|^{4}<\infty,$
$\sup_{N\geq 1}\mathbf{E}|\Phi^{N}(t)-\Phi^{N}(s)|^{2}+\mathbf{E}|\Phi(t)-\Phi(s)|^{2}\leq K|t-s|.$
@@
$V(t)=\int_{0}^{t}\nabla\bar{f}(Y(s))V(s)\mathrm{d}s+\int_{0}^{t}\nabla g(Y(s))V(s)\mathrm{d}W(s)+\sum_{i=1}^{4}I_{i}(t),~t\in[0,T].$